OpenStudy (anonymous):

Which are real zeroes of this function? x3 + 2x2 – 9x – 18

OpenStudy (mathmath333):

hint : try to factor \(x^2\) out

OpenStudy (anonymous):



OpenStudy (anonymous):

@satellite73

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{hint:\\~\\x^3 + 2x^2 – 9x – 18 \hspace{.33em}\\~\\ =x^2(x + 2)– 9(x + 2) \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

What do I do with that hint?

OpenStudy (mathmath333):

good question

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{x^3 + 2x^2 – 9x – 18 \hspace{.33em}\\~\\ =x^2(x + 2)– 9(x + 2) \hspace{.33em}\\~\\ =(x^2-9)(x + 2) \hspace{.33em}\\~\\ =(x^2-3^2)(x + 2) \hspace{.33em}\\~\\ =(x-3)(x+3)(x + 2) \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

still need help

OpenStudy (anonymous):

x=-3,3,2 are the zeroes

OpenStudy (anonymous):

THank you. So the ral zero's of the function are 3, -3, 2.

OpenStudy (anonymous):

real*

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{(x-3)(x+3)(x + 2) =0 \hspace{.33em}\\~\\ }\end{align}\) find 3 x values u will get the same.

OpenStudy (anonymous):

THank you Grandmaster.